how do I solve 9(19x-4)=3x
Answer by Alan3354(6097)

how do I solve 9(19x-4)=3x
----------
9(19x-4)=3x
Get rid of the parentheses
171x - 36 = 3x
Get all the x terms on one side. Subtract 3x from both sides
168x - 36 = 0
Add 36
168x = 36
Divide by 168
x = 36/168
x = 3/14